what is difference between B and H?


  1. what is difference between B and H?

Grade:9

2 Answers

Dastagir Hassain
11 years ago

H IS  INTENSITY OF MAGNETIC FIELD

B IS MAGNETIC FIELD INDUCTION

 

H=B IN C.G.S SYSTEM

BUT....,

B=uoH

H= m / (4pi r^2)

H =B/uo

yaswanth kumar madugula
11 years ago

hai sr i know very little much sir.

 

Based on Maxwell''s equations, electric fields are generated by changing B fields, while H fields are generated by changing electric fields.
In dc fields, static electric E fields create currents (magnetization currents) I (when σ>0) which in turn produce static H fields. A static B field (actually d/dt ∫B·n dA = 0) cannot produce an electric field E.

Even if we used natural units where μ0 = ε0 = 1, this distinction between B and H remains. In magnetic materials, B is not linearly related to H due to the magnetization term M. If B = H in space, then BH in magnetic materials.